# Energy-Efficiency and Spectral-Efficiency Trade-off in Distributed Massive-MIMO Networks

### v2

@@ I Introduction
In the era of 5G and beyond, the fast expansion of connected devices and exponential rise in data traffic have made spectral efficiency (SE) and energy efficiency (EE) key network design factors. Enhanced SE is essential to meet the growing demand for higher data rates and better quality of service, but it often comes at the expense of significantly increased energy consumption due to higher power demands on infrastructure and devices. As the world moves toward green communication to reduce environmental effect, thus establishing the trade-off between the EE and SE becomes a critical challenge in 5G networks, where technologies such as massive MIMO and dense deployment of access points aim to boost SE but risk increasing energy consumption to unsustainable levels, both environmentally and economically. Therefore, striking a balance between SE and EE is not only a technical necessity, but also a critical step towards achieving economically and environmentally sustainable communication systems.
@@
Distributed massive-MIMO (D-mMIMO) combines the advantages of massive-MIMO and the cell-free architecture to improve performance of the network [1]. D-mMIMO systems provide uniform service quality over a large coverage area by distributing antennas across the region, significantly enhancing both EE and SE compared to conventional cellular networks [2].
Interestingly, in general, there is an inherent trade-off between the EE and SE, increasing the SE may eventually lead to a decease in EE, and vice versa. This trade-off can also be observed in D-mMIMO systems[3]. In [3], the authors have analysed the EE-SE trade-off for downlink scenario, but these EE and SE are not optimized. Power allocation, AP-UE association, and many other factors impact both, the EE and SE of the D-mMIMO[2]. Thus, there is a need to analyse the EE-SE trade-off after performing the optimal resource allocation.
@@
One of the primary factors that influence both EE and SE is power allocation. Another important aspect in D-mMIMO systems is the AP-UE association, where each user is dynamically associated with a subset of APs that collaboratively serve the user. This association plays a vital role in both the EE and SE. Thus, the joint power allocation and AP-UE association is important to enhance the energy and spectral efficiencies of D-mMIMO networks and to analyze their trade-off.
@@
Much work has been done for power allocation and AP-UE association to enhance either SE or EE of the D-mMIMO systems [2, 4, 5, 6, 7, 8, 9, 10, 11]. But none of the existing work, to the best of our knowledge, deals with the EE-SE trade-off while performing AP-UE association and power allocation for the D-mMIMO.
@@
In this paper, we formulate an optimization problem for EE maximization with a constraint on the quality-of-service (QoS) requirement of the sum SE with respect to power allocation and AP-UE association. The goal is to establish a trade-off between EE and SE while finding the power allocation across APs and the association between APs and UEs that maximize network performance under these constraints.
@@
Organization: The paper is organized as follows: Section II presents the system model highlighting the key components and assumptions used in this study. Section III details the problem formulation. Section IV introduces our proposed solution and establishes its convergence. In Section V, using the proposed solution approach we numerically evaluate the trade-off between the EE and SE with respect to various system parameters. Finally, Section VI concludes our work and provides some directions for future work.
@@ II System Model
We consider a scenario where $T$ single antenna UEs and $M$ multi-antenna APs operate within a defined coverage area. Each AP is equipped with $A$ antennas, and the relationship $T\ll MA$ indicates a high antenna-to-user ratio, facilitating substantial spatial multiplexing gains. The APs and the UEs are uniformly distributed across the coverage area, ensuring that every UE is within the service range of multiple APs. All APs can serve all UEs simultaneously using the same time and frequency resources. Connectivity between the APs and the central processing unit (CPU) is ensured via a front-haul link, which is critical for coordinating signal processing tasks and distributing user data and control information efficiently across the network as shown in Fig. 1. To ensure network scalability, a user-centric approach has been adopted, where each UE is served by a subset of APs, as depicted in Fig. 1.
Figure 1: A typical network model. Here each oval identifies the UE(s) and the APs that serve it(them).
This setup uses Time-Division Duplexing (TDD) mode for communication. Channel estimation is conducted during the uplink phase where the UEs transmit orthogonal pilot signals, and the APs estimate the uplink channel vectors. The system employs a block fading model, characterized by a coherence block of length $L_{c}$. Within each coherence block, a portion defined by $L_{p}$ is dedicated to pilot transmission for channel estimation. The small-scale fading component, denoted by $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, follows a Rayleigh distribution. Each component of $\textbf{h}_{mt}$ is assumed to be an independent and identically distributed (i.i.d.) complex Gaussian random variable with zero mean and unit variance. Large-scale fading coefficient (LSFC), denoted by $\beta_{mt}$, accounts for path-loss and shadowing effects. The overall channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ is given by the product $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
@@ II-A Uplink Pilot Training
$\textbf{y}^{pilot}_{m}=\sum_{t=1}^{T}\sqrt{L_{p}p_{p}}\textbf{g}_{mt}\mathbf{\Psi}_{t}+\textbf{n}_{m},$
where $\textbf{n}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ is the complex additive white Gaussian noise matrix with each element distributed normally with zero mean and unit variance. Also, $p_{p}$ represents the normalized signal-to-noise power for pilot transmission. The MMSE estimate $\hat{\textbf{g}}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$ of $\textbf{g}_{mt}$ and the mean-square of the estimates are given by [2]:
$\hat{\textbf{g}}_{mt}=\frac{\sqrt{L_{p}p_{p}}\beta_{mk}}{\sum_{t^{\prime}=1}^{T}L_{p}p_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t},$
$\gamma_{mt}=\frac{L_{p}p_{p}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}L_{p}p_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}.$
@@ II-B Uplink Data Transmission
where $x_{t}$ is the uplink payload signal transmitted from the UE $t$, such that $\mathbb{E}\{|x_{t}|^{2}\}=1$. The uplink normalized signal-to-noise power and power control coefficient of the UE $t$ are denoted by $p_{u}$ and $\eta^{u}$, respectively. The large scale fading decoding (LSFD) of the UE $t$ for all APs is performed at CPU after local data estimation using the combining vector, $\textbf{v}_{mt}$ for APs $m$, as :
$\hat{y}_{t}=\sum_{m=1}^{M}d_{mt}\hat{y}_{mt}=\sum_{m=1}^{M}d_{mt}\textbf{v}_{mt}^{H}\textbf{y}^{u}_{m},$ (1)
where $d_{mt}$ is the AP-UE association binary variable of the UE $t$ for AP $m$ such that when AP $m$ serves the UE $t$, then $d_{mt}=1$, else $d_{mt}=0$. Eq. (1) can be re-written as :
$\begin{split}\hat{y}_{t}&={\sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}d_{mt}\mathbb{E}\left\{\textbf{v}_{mt}^{H}\textbf{g}_{mt}\right\}x_{t}}+{\sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}d_{mt}(\textbf{v}_{mt}^{H}\textbf{g}_{mt}-\mathbb{E}\left\{\textbf{v}_{mt}^{H}\textbf{g}_{mt}\right\})x_{t}}\\ &+{\sum_{k\in\mathcal{P}_{t}/\{t\}}\ \sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}d_{mt}\textbf{v}_{mt}^{H}\textbf{g}_{mk}x_{k}}+{\sum_{k\notin\mathcal{P}_{t}}\ \sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}d_{mt}\textbf{v}_{mt}^{H}\textbf{g}_{mk}x_{k}}+{\sum_{m=1}^{M}d_{mt}\textbf{v}_{mt}^{H}\textbf{n}_{m}},\end{split}$
@@ II-C Spectral Efficiency and Energy Efficiency
The uplink spectral efficiency can be given by :
$\text{SE}^{u}_{t}=w\log_{2}(1+\Gamma_{t}),$ (2)
where $w=(1-\frac{L_{p}}{L_{c}})$ and $\Gamma_{t}$ is the effective uplink signal-to-interference-noise ratio (SINR). The closed-form expression of effective SINR for any UE $t$ the partial full-pilot zero-forcing (PFZF) combining[12] is given by :
$\Gamma_{t}=\frac{\eta^{u}_{t}p_{u}\left|\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}\gamma_{mt}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}\gamma_{mt}\right|^{2}}{I_{t}},$ (3)
where
$\begin{split}I_{t}&=\text{A}_{t}^{\text{P}}+\text{B}_{t}^{\text{P}}+\text{C}_{t}^{\text{P}},\\ \text{A}_{t}^{\text{P}}&=\sum_{t^{\prime}=1}^{T}\eta^{u}_{t^{\prime}}p_{u}\Bigg{(}\sum\limits_{\ \ m\in\mathcal{Z}_{t}}\frac{d_{mt}^{2}\gamma_{mt}(\beta_{mt^{\prime}}{-}\gamma_{mt^{\prime}})}{A-L_{\mathcal{S}_{m}}}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}^{2}\gamma_{mt}\beta_{mt^{\prime}}\!\!\Bigg{)},\\ \text{B}_{t}^{\text{P}}&=\sum_{\ \ t^{\prime}\in\mathcal{P}_{t}/\{t\}}\eta^{u}_{t^{\prime}}p_{u}\Bigg{|}\sum\limits_{\ \ m\in\mathcal{Z}_{t}}\frac{d_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}+\!A\sum\limits_{m\in\mathcal{Q}_{t}}\frac{d_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}\Bigg{|}^{2},\\ \text{C}_{t}^{\text{P}}&=\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}^{2}\frac{\gamma_{mt}}{A-L_{\mathcal{S}_{m}}}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}^{2}\gamma_{mt}.\end{split}$
The total uplink energy efficiency of the network in bits per joule is defined as:
$\text{EE}=\frac{wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P_{\text{T}}},$ (4)
where $B$ is the total bandwidth and according to [13],
$P_{\text{T}}=P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+wP^{\text{deco}}_{cpu}B\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t}),$
${P}^{c}_{\text{T}}=\sum_{t=1}^{T}\Bigg{(}\Big{(}\!\frac{\eta^{u}_{t}p_{n}p_{u}}{\zeta}+d_{mt}P^{\text{lsfd}}_{cpu}\!\Big{)}+\sum_{m=1}^{M}\Big{(}Ad_{mt}P^{\text{proc}}+d_{mt}P^{\text{sig}}\Big{)}\Bigg{)},$
@@ III Problem Formulation
In this section, we introduce an optimization problem formulation for maximizing EE that involves allocating power to all UEs and determining AP-UE allocations, subject to the constraint of achieving a minimum total SE. The problem formulated is given by:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}B\sum_{t=1}^{T}\text{SE}^{u}_{t}},$ (5a)
@@
$w\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (5d)
@@
where $\bm{\eta^{u}}=\{\eta^{u}_{t}\}_{t\in\mathcal{T}}$ and D is the AP-UE association matrix with each element as $d_{mt}$. $\textbf{D}_{t}$ represents the column vector of matrix D with respect to the UE $t$. The regularization coefficient $\alpha\geq 0$ is the trade-off factor between the front-haul load and the SE. $\text{SE}^{\text{QoS}}$ represents the minimum QoS requirement. Constraint (5b) represents the binary variable indicating the association between the UEs and APs. Constraint (5c) represents the uplink power control coefficient. Constraint (5d) is for the minimum QoS requirement and constraint (5e) specifies that at least one AP must be connected to the UE. The above Problem (5) is an instance of mixed-integer non-linear programming (MINLP) problems which are, generally, NP-hard. Thus, to find a solution with lower complexity, we employ the Successive Convex Approximation (SCA) technique[14], as described next.

### v3

@@ I Introduction
In the era of 5G and beyond, the fast expansion of connected devices and exponential rise in data traffic have made spectral efficiency (SE) and energy efficiency (EE) key network design factors. Enhanced SE is desirable to meet the growing demand for higher data rates and better quality of service, but it often comes at the cost of significantly increased energy consumption due to higher power demands on infrastructure and devices. As we increasingly move toward green communication, optimizing the EE and SE becomes a critical challenge in 5G networks, where technologies such as massive MIMO and dense deployment of access points aim to boost SE, but may increase energy consumption to unsustainable levels, both environmentally and economically. Therefore, striking a balance between SE and EE is not only desirable technically, but is also a critical step towards sustainable communication systems.
@@
Distributed massive-MIMO (D-mMIMO) combines the advantages of massive-MIMO and the cell-free architecture to improve performance of the network [1]. The D-mMIMO systems provide uniform service quality over a coverage area by distributing antennas across the region, enhancing both EE and SE compared to conventional cellular networks [2].
@@
In general, there is an inherent trade-off between the EE and SE, increasing the SE may lead to a decrease in the EE, and vice versa. This trade-off can also be observed in D-mMIMO systems. In [3], the authors have analyzed the EE-SE trade-off for downlink scenarios, but the EE and SE are not optimized. In D-mMIMO systems, two of the primary factors that influence both EE and SE are power allocation and AP-UE association, where each user is dynamically associated with a subset of APs that collaboratively serve the user [2]. Thus, the joint power allocation and AP-UE association is important to enhance the energy and spectral efficiencies of D-mMIMO networks and to analyze their trade-off.
@@
Much work has been done for power allocation and AP-UE association to enhance either SE or EE of the D-mMIMO systems [2, 4, 5, 6, 7, 8, 9, 10, 11]. But none of the existing work, to the best of our knowledge, deals with the EE-SE trade-off while performing AP-UE association and power allocation for the D-mMIMO. Our work attempts to address this gap.
@@
We formulate an optimization problem for EE maximization with a constraint on the quality-of-service (QoS) requirement of the sum SE with respect to power allocation and AP-UE association. The goal is to establish a trade-off between EE and SE while finding the power allocation across APs and the association between APs and UEs that maximize network performance under these constraints.
@@
Organization: Section II presents the system model. Section III details the problem formulation. Section IV introduces our proposed solution and establishes its convergence. In Section V, using the proposed solution approach, we numerically evaluate the trade-off between the EE and SE with respect to various system parameters. Finally, Section VI concludes our work and provides some directions for future work.
@@ II System Model
We consider a scenario where $T$ single antenna UEs and $M$ multi-antenna APs operate within a defined coverage area. Each AP is equipped with $A$ antennas, and the relationship $T\ll MA$ indicates a high antenna-to-user ratio, facilitating substantial spatial multiplexing gains. The APs and the UEs are uniformly distributed across the coverage area, ensuring that every UE is within the service range of multiple APs. All APs can serve all UEs simultaneously using the same time and frequency resources. Connectivity between the APs and the central processing unit (CPU) is ensured via a front-haul link, which is critical for coordinating signal processing tasks and distributing user data and control information efficiently across the network. To ensure network scalability, a user-centric approach has been adopted, where each UE is served by a subset of APs, as depicted in Fig. 1.
Figure 1: A typical network model. Here each colored oval identifies the UE(s) and the APs that serve it(them).
This setup uses Time-Division Duplexing (TDD) mode for communication. Channel estimation is conducted during the uplink phase where the UEs transmit orthogonal pilot signals, and the APs estimate the uplink channel vectors. The system employs a block fading model, characterized by a coherence block of length $L_{c}$. Within each block, a portion defined by $L_{p}$ is dedicated to pilot transmission for channel estimation. The small-scale fading component, denoted by $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, follows a Rayleigh distribution. Each component of $\textbf{h}_{mt}$ is assumed to be an independent and identically distributed (i.i.d.) complex Gaussian random variable with zero mean and unit variance. Large-scale fading coefficient (LSFC), denoted by $\beta_{mt}$, accounts for path-loss and shadowing effects. The overall channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ is given by $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
@@ II-A Uplink Pilot Training
$\textbf{y}^{pilot}_{m}=\sum_{t=1}^{T}\sqrt{L_{p}p_{p}}\textbf{g}_{mt}\mathbf{\Psi}^{\text{H}}_{t}+\textbf{N}_{m},$
where $\textbf{N}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ is the complex additive white Gaussian noise matrix with each element distributed normally with zero mean and unit variance. Also, $p_{p}$ represents the normalized signal-to-noise power for pilot transmission. The mean square value of any component of MMSE-estimated channel vector $\textbf{g}_{mt}$ is given by [2]: $\gamma_{mt}=\dfrac{L_{p}p_{p}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}L_{p}p_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}$.
@@ II-B Uplink Data Transmission
where $x_{t}$ is the uplink payload signal transmitted from the UE $t$, such that $\mathbb{E}\{|x_{t}|^{2}\}=1$. The uplink normalized signal-to-noise power and power control coefficient of the UE $t$ are denoted by $p_{u}$ and $\eta^{u}$, respectively. Also, $\textbf{n}_{m}$ is the complex additive white Gaussian noise matrix with each element distributed normally with zero mean and unit variance. The large scale fading decoding (LSFD) of the UE $t$ for all APs is performed at CPU after local data estimation using the combining vector, $\textbf{v}_{mt}$ for APs $m$, as: $\hat{y}_{t}=\sum_{m=1}^{M}d_{mt}\textbf{v}_{mt}^{H}\textbf{y}^{u}_{m}$, where $d_{mt}$ is the AP-UE association binary variable of the UE $t$ for AP $m$ such that when AP $m$ serves the UE $t$, then $d_{mt}=1$, else $d_{mt}=0$.
@@ II-C Spectral Efficiency and Energy Efficiency
The uplink spectral efficiency of the UE $t$ and of the overall system are given by:
$\text{SE}^{u}_{t}$ $=w\log_{2}(1+\Gamma_{t}),$
$\text{SE}^{u}|_{\Gamma_{t}}$ $=\sum_{t=1}^{T}\text{SE}^{u}_{t},$
where $w=(1-\frac{L_{p}}{L_{c}})$ and $\Gamma_{t}$ is the effective uplink signal-to-interference-noise ratio (SINR) of the UE $t$.
The closed-form expression of effective SINR for any UE $t$ the partial full-pilot zero-forcing (PFZF) combining[12] is given by $\Gamma_{t}=\text{DS}_{t}/I_{t}$, where
$\begin{split}\text{DS}_{t}&=\eta^{u}_{t}p_{u}\Big{|}\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}\gamma_{mt}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}\gamma_{mt}\Big{|}^{2},\\ I_{t}&=\text{A}_{t}^{\text{P}}+\text{B}_{t}^{\text{P}}+\text{C}_{t}^{\text{P}},\\ \text{A}_{t}^{\text{P}}&=\sum_{t^{\prime}=1}^{T}\!\!\eta^{u}_{t^{\prime}}p_{u}\Big{(}\!\!\!\!\!\sum\limits_{m\in\mathcal{Z}_{t}}\frac{d_{mt}^{2}\gamma_{mt}(\beta_{mt^{\prime}}{-}\gamma_{mt^{\prime}})}{A-L_{\mathcal{S}_{m}}}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}^{2}\gamma_{mt}\beta_{mt^{\prime}}\Big{)}\\ \text{B}_{t}^{\text{P}}&=\sum_{t^{\prime}\in\mathcal{P}_{t}/\{t\}}\eta^{u}_{t^{\prime}}p_{u}\Big{|}\sum\limits_{m\in\mathcal{Z}_{t}}\frac{d_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}+\!A\sum\limits_{m\in\mathcal{Q}_{t}}\frac{d_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}\!\Big{|}^{2}\\ \text{C}_{t}^{\text{P}}&=\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}^{2}\frac{\gamma_{mt}}{A-L_{\mathcal{S}_{m}}}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}^{2}\gamma_{mt}.\end{split}$
The total uplink energy efficiency of the network in bits per joule is defined as: $\text{EE}=\dfrac{B\,\text{SE}^{u}|_{\Gamma_{t}}}{P_{\text{T}}}$, where $B$ is the total bandwidth and according to [13]:
$P_{\text{T}}=P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}B\,\text{SE}^{u}|_{\Gamma_{t}},$
${P}^{c}_{\text{T}}=\sum_{t=1}^{T}\Big{(}\frac{\eta^{u}_{t}p_{n}p_{u}}{\zeta}+\sum_{m=1}^{M}\big{(}d_{mt}P^{\text{lsfd}}_{cpu}\!+Ad_{mt}P^{\text{proc}}+d_{mt}P^{\text{sig}}\big{)}\Big{)},$
@@ III Problem Formulation
We introduce an optimization problem formulation for maximizing EE that involves allocating power to all UEs and determining AP-UE allocations, subject to the constraint of achieving a minimum total SE as follows.
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{B\,\text{SE}^{u}|_{\Gamma_{t}}}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}B\,\text{SE}^{u}|_{\Gamma_{t}}},$ (1a)
@@
$\text{SE}^{u}|_{\Gamma_{t}}\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (1d)
@@
where $\bm{\eta^{u}}=\{\eta^{u}_{t}\}_{t\in\mathcal{T}}$ and D is the AP-UE association matrix with each element as $d_{mt}$. $\textbf{D}_{t}$ represents the column vector of matrix D with respect to the UE $t$. The regularization coefficient $\alpha\geq 0$ is the trade-off factor between the front-haul load and the SE. $\text{SE}^{\text{QoS}}$ represents the minimum QoS requirement. Constraint (1b) represents the binary variable indicating the association between the UEs and APs. Constraint (1c) represents the uplink power control coefficient. Constraint (1d) is for the minimum QoS requirement and constraint (1e) specifies that at least one AP must be connected to the UE.
This problem is an instance of mixed-integer non-linear programming (MINLP) problems which are, generally, NP-hard. We employ the Successive Convex Approximation (SCA) technique[14] to find an approximate but lower complexity solution for it, as described next.